Could you please answer 2 + 3 as they were wrong in the last post? Inventory Costing Methods Carrington Inc. reported the following information for the month of February: Inventory, February 1
Answer #1
Unit Unit cost Total Cost
Beg Inv 63 26 1638
Purchase 1 55 27 1485
Purchase 2 60 29 1740
Purchase 3 45 31 1395
Total 223 6258

Calculate following

Assumption Cost of goods sold Ending inventory
1 Of the 142 units sold, 53 Cost $26, 39 cost $27, 45 Cost $29 and 5 Cost $31 3891 6258-3891 = 2367
2 FIFO 6258-2439 = 3819 1395+36*29 = 2439
3 LIFO 6258-2124 = 4134 1638+18*27 = 2124
4 Weighted average 6258-2273 = 3985 81*28.06 = 2273